n0mn0mlemon:

Which equations support the fact that rational numbers are closed under subtraction? Select each correct answer. 4.5−0.5=4 2√3−√3=√3 √8−√8=0 5√4−√4=4√4 Please help!

n0mn0mlemon:

I don't know, that's the question I was asked and that's why I need help.

n0mn0mlemon:

I don't know what a snap pic is

jimthompson5910:

If you take any two rational numbers, subtract them, and get a rational number as a result, then this means the set of rational number is closed under subtraction. In other words, let A and B be rational numbers C = A-B is also rational if rational numbers was closed under subtraction

jimthompson5910:

For choice A we have 4.5 = 9/2 which is rational 0.5 = 1/2 also rational 4.5−0.5=4 which is rational because 4 = 4/1 So this is one example where subtracting two rational numbers leads to some other rational number. It shows closure. There's one other answer.

n0mn0mlemon:

What is it?

jimthompson5910:

Note how \(\sqrt{4} = 2 = \frac{2}{1}\) which shows it's rational This means \(5\sqrt{4}\) and \(4\sqrt{4}\) are also rational. This is sufficient to show that D is the other answer that shows an example of closure under subtraction.

jimthompson5910:

Another way to phrase closure is to say (rational) - (rational) = rational

n0mn0mlemon:

Ohhh, okay. Thank you!

jimthompson5910:

No problem